# Reconstructing Quantitative Cerebral Perfusion Images Directly From Measured Sinogram Data Acquired Using C-arm Cone-Beam CT

### v1

@@ II-E The Joint Optimization Problem in TRAINER
where $\mathbf{V}$ denotes the dual variable of $\mathbf{U}$ and $\lambda$ controls the tradeoff between data fitting term and the regularization term.
The unconstrained problem can be efficiently solved by the Alternating Direction Method of Multipliers (ADMM) algorithm iteratively in the following three sub-problems:
@@ II-F Numerical Optimization Algorithm
The optimization of sub-problem Eq. (15) can be technically challenging. As we can observe, the cost function in Eq. (15) is highly nonlinear and non-convex with respect to the unknown perfusion parametric images $\mathbf{U}$. Here, we choose to use the Adam optimizer [29] to minimize the cost function defined in Eq. (15). The Adam optimizer is particularly suitable for this task due to its ability to handle non-smooth objective functions and dynamically adapt learning rates to facilitate faster convergence.
@@
Sub-problem Eq. (17) is a nonlinear least square problem and it has the same form as the conventional supervised training problem, using $\mathbf{U}^{(k+1)}+\mathbf{V}^{(k)}$ as the training labels. The gradient descent has been employed to solve this supervised training problem. Following Eq. (17), we regularize the problem by considering the re-parametrization $\mathcal{G}_{\mathbf{\Theta}}$ and optimizing the resulting energy w.r.t. $\mathbf{\Theta}$. Optimization still applies gradient descent as the optimizer, exploiting the fact that both the neural network and the most common downsampling operators, such as Lanczos, are differentiable [30].
@@ II-F3 Pseudo-code to Implement TRAINER
To implement TRAINER, we first need to initialize $\mathbf{U}$, $\mathbf{V}$, and $\mathbf{Z}$, and then optimize these arguments iteratively. The perfusion parametric reconstruction sub-problem (Eq. (15)) has been solved by a iterative process typically performed for 300 iterations to obtain $\mathbf{U}^{(k+1)}$. The conditional generative model training sub-problem (Eq. (17)) has been solved by a iterative process typically performed for 2000 iterations to obtain $\mathbf{\Theta}^{(k+1)}$. Finally, the Lagrangian multiplier $\mathbf{V}$ is updated, and this iterative process continues until convergence criteria are met.
@@
1:  Input: line integral data $\vec{\mathbf{Y}}$, Initialized $\mathbf{CBF}$ and $\mathbf{MTT}$
2:  Output: $\mathbf{U}=\{\mathbf{CBF},\mathbf{MTT}\}$
3:  Initialize: $\mathbf{V}^{(0)},\mathbf{\Theta}^{(0)},\mathbf{Z},\lambda,\alpha,\eta$
4:  while $\mathcal{L}_{\lambda}(\mathbf{U}^{(k)},\mathbf{\Theta}^{(k)},\mathbf{V}^{(k)})\geq\varepsilon$ do
5:     while $k\leq\text{Num Of ADMM}$ do
6:        $q=0$
7:        while $q<\mathbf{Q}$ do
8:           $\mathbf{U}^{(q+1)}=\mathbf{U}^{(q)}-\eta\nabla_{\mathbf{U}}\bigg{[}\frac{1}{2}||\vec{\mathbf{Y}}-\overrightarrow{\bar{\mathbf{Y}}(\mathbf{U})}||_{\mathbf{W}}^{2}+\frac{\lambda}{2}||\mathbf{U}-\mathcal{G}_{\mathbf{\Theta}^{(k)}}(\mathbf{Z})+\mathbf{V}^{(k)}||^{2}\bigg{]}$
[…]
10:        end while
11:        $\mathbf{U}^{(k+1)}=\mathbf{U}^{(\mathbf{Q})}$
12:        $p=0$
13:        while $p<\mathbf{P}$ do
14:           $\mathbf{\Theta}^{(p+1)}=\mathbf{\Theta}^{(p)}-\alpha\nabla_{\mathbf{\Theta}}||\mathbf{U}^{(k+1)}+\mathbf{V}^{(k)}-\mathcal{G}_{\mathbf{\Theta}}(\mathbf{Z})||^{2}$
[…]
16:        end while
17:        $\mathcal{G}_{\mathbf{\Theta}^{(k+1)}}(\mathbf{Z})=\mathcal{G}_{\mathbf{\Theta^{(\mathbf{P})}}}(\mathbf{Z})$
18:        $\mathbf{V}^{(k+1)}=\mathbf{V}^{(k)}+\mathbf{U}^{(k+1)}-\mathcal{G}_{\mathbf{\Theta}^{(k+1)}}(\mathbf{Z})$
[…]
20:     end while
@@ II-F4 Network Architecture
The architecture of $\mathcal{G}_{\mathbf{\Theta}}$ is illustrated in the Fig. 1 (a). The architecture is designed based on an encoder-decoder structure with skip connections. Parameters $n_{u}[i]$, $n_{d}[i]$, and $n_{s}[i]$ denote the number of filters for upsampling, downsampling, and skip connections at the $i$-th depth, respectively. Parameters $k_{u}[i]$, $k_{d}[i]$, and $k_{s}[i]$ denote the kernel sizes at the $i$-th depth respectively [30]. Bilinear interpolation is employed for the upsampling process. We use LeakyReLU [31] to introduce the non-linearity. During the iterative training process of network parameters $\mathbf{\Theta}$, we add additive normal noise with a mean of zero values and a standard deviation of $\sigma_{p}=\frac{1}{30}$ to input $\mathbf{Z}$ for perturbation. Experimental results show that this method yields superior performance. Additionally, previous research indicates that the optimization process tends to become unstable as the loss decreases and approaches a specific threshold. This instability is characterized by a notable increase in the loss value and the appearance of blurriness in the generated image $\mathcal{G}_{\mathbf{\Theta}}$. To make the training process more stable, we revert the parameters to those from the previous iteration if the loss difference between two consecutive iterations exceeds a specified threshold [30].
Figure 1: (a) The architecture of $\mathcal{G}$. $n_{u}[i]$, $n_{d}[i]$, and $n_{s}[i]$ denote the number of filters for upsampling, downsampling, and skip connections at the $i$-th depth $i$ respectively. Parameters $k_{u}[i]$, $k_{d}[i]$, and $k_{s}[i]$ denote the kernel sizes at the $i$-th depth $i$ respectively. (b) Conditional image generation. During each iteration, the network parameters $\mathbf{\Theta}$ determine the image $\mathcal{G}_{\mathbf{\Theta}}(\mathbf{Z})$, and the mapping $\mathcal{G}$ is represented as a deep convolutional neural network parameterized by $\mathbf{\Theta}$.
@@ II-G Implementation Details
II-G1 Obtaining AIF
To apply TRAINER, it is necessary to accurately reconstruct the AIF and then reconstruct the perfusion parametric images using the AIF obtained. In numerical simulation studies, we deliberately applied the same AIF across all baseline methods and TRAINER to establish fair comparisons. By decoupling the AIF reconstruction from the perfusion parametric image reconstruction, the contributions of different parametric image reconstruction approaches can be highlighted. In numerical simulation studies, we applied the ground truth AIF to generate the corresponding parametric images using all baseline methods and TRAINER.
II-G2 Initialization and Hyperparameter Setting
We first initialized the iterative sequence via the following empirical schemes to obtained the warm initialization of $\mathbf{U}$. The value of CBF for each pixel is initialized as 40 mL/100g/min and MTT for each pixel is initialized as 10 seconds because such initialization helps the iterative process converge more quickly to the expected solution. $\mathbf{\Theta}$ is initialized as random values uniformly distributed between 0 and 1, while $\mathbf{V}$ is initialized as 0. The variable $\mathbf{Z}$ is initialized within the interval $[0,1/10)$, following a uniform distribution. To optimize the performance of TRAINER, the hyperparameters $\lambda$, $\alpha,\eta$ for numerical simulation studies are empirically set to 0.0015, 0.01, and 0.001, respectively.
II-G3 Other Remarks
@@ III-A Numerical Simulation Studies with Known Ground Truth
The design of the numerical phantom in this study was based on a digital anthropomorphic perfusion phantom [32]. Regions corresponding to artery, healthy gray matter, healthy white matter, cerebrospinal fluid, penumbra and ischemic core are defined in the phantom design software toolkit. The residual functions are defined according to programmed cerebral blood flow and mean transit time values. The ground true TAC of each pixel is then calculated according to the indicator dilution theory. The data acquisition geometry was designed to simulate the scans performed in the animal subject study.
@@
Ideal projection data were generated using a ray-driven implementation of fan-beam forward projection. Projections with quantum noise were then produced by sampling from a Poisson distribution for each projection ray of the ideal data. For CBCT perfusion data acquisitions, entrance photon numbers $I_{0}=1\times 10^{6}$ photons per ray were used, such that, noise levels were deliberately selected to match the image noise measured for standard exposure level. We denote this exposure level as the full exposure level. Entrance photon number per ray $I_{0}=1\times 10^{5},5\times 10^{4},1\times 10^{4}$ were used to simulated lower exposure levels respectively.
@@ III-B Baseline Methods for Performance Comparison
Four baseline methods are compared: Singular value decomposition with Tikhonov regularization [33], and Tensor Total-Variation Regularization (TTV) [34], temporal recovery technique (TRT) [10], and eSMART-RECON [6]. For the SVD-based and Tikhonov algorithms, threshold value of 0.15 (15% of the maximum singular value) is empirically chosen to achieve optimal performance. These methods are widely recognized as regularized deconvolution techniques for CT perfusion imaging and are extensively incorporated into commercial medical software [35]. TTV [34] applies total variation regularization to the impulse response function, integrating spatial vascular structure correlation and temporal blood signal continuity. This algorithm converges rapidly and exhibits low computational complexity. TRT algorithm introduces a novel temporal recovery technique designed to restore the contrast-enhanced curves in C-arm CBCT perfusion imaging, addressing the limitations of existing C-arm CBCT systems in terms of data acquisition speed and temporal sampling density. eSMART improves the temporal resolution of C-arm CBCT via incorporating the prior knowledge of periodicity of the limited-view artifacts in the C-arm CBCT perfusion data acquisition scheme. To improve image quality and quantitative accuracy at lower exposure levels, noise reduction in the time-resolved images reconstructed by each baseline method was achieved through spatial and temporal domain filtering. Specifically, spatial filtering was performed using a Gaussian filter with a standard deviation of 1 and a kernel size of $5\times 5$, while temporal filtering was conducted with a Gaussian filter having a standard deviation of 1 and a kernel size of $3\times 1$. Perfusion parametric images of each baseline method were then generated using the correspondingly denoised time-resolved images.
@@ III-C Metrics for Quantitative Performance Assessment
The study employs normalized root mean square error (nrMSE) and Peak Signal-to-Noise Ratio (PSNR) as evaluation metrics.
$\mathrm{nrMSE}=\frac{\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-y_{i}^{truth}\right)^{2}}}{\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}^{truth})^{2}}}\times 100\%,$ (20)
where $y$ denotes the image being assessed, $y^{\text{truth}}$ denotes the ground truth image, and the subscript $i$ denotes the $i$-th voxel index of the image.
@@ IV-B Qualitative Assessment of Perfusion Parametric Images
Quantitative perfusion images reconstructed at the gantry rotation speed of $dt=8$ sec and four exposure levels are shown in Fig. 3 and Fig. 4. For CBF images, at the full exposure level, compared with the ground truth, FBP+Tikh and FBP+TTV underestimate healthy tissues while overestimate diseased tissues (penumbra and ischemic core). TRT+TTV underestimates but eSMART+TTV overestimates healthy white matters. In addition, both TRT+TTV and eSMART+TTV overestimate diseased tissues. In contrast, TRAINER can accurately reconstruct CBF values for all types of tissues. When the exposure level is getting lower, penumbra and ischemic core are not be distinguished in baseline methods. Error and noise level of CBF images reconstructed by baseline methods are getting larger. For MTT images, at the full exposure level, compared with the ground truth, baseline methods underestimate all tissue types. Penumbra and ischemic core are not be distinguished in baseline methods even at the full exposure level. TRAINER accurately estimate the value of all types of tissues. When the exposure level is getting lower, error and noise level of MTT images reconstructed by baseline methods are getting larger. Results shown in Fig. 3 and Fig. 4 demonstrated that TRAINER can accurately reconstruct quantitative perfusion images at gantry rotation speed of $dt=8$ sec even at the lowest exposure level for all tissue types while other methods cannot.
@@
To validate the effectiveness of the proposed TRAINER and also explain why other baseline methods cannot obtain accurate quantitative perfusion images, we presented the time-activity curves (TACs) of the reconstructed time-resolved images using TRAINER and compared them to those obtained from the baseline methods.
As shown in Fig. 5, TRAINER consistently generates TACs closer to the ground truth (Truth) across three different tissue types: healthy tissue, penumbra, and core. Specifically, in the healthy tissue region, TACs obtained using TRAINER are nearly consistent to Truth, demonstrating a significant improvement on reconstruction accuracy over the baseline methods. In the penumbra region, although all methods show some fluctuations, TACs generated by TRAINER remains closer to Truth. In the core region, TRAINER yields smoother and more accurate TACs compared to the baseline methods, further highlighting the superior quantitative accuracy of TRAINER. These results further confirm the efficacy of TRAINER, demonstrating its superior performance in reconstructing TACs using data acquired by slow rotation speed systems.
Figure 5: Comparison of time-activity curves for different tissue types. The plots illustrate the time-activity curves for healthy tissue (top), penumbra (middle), and ischemic core (bottom). The TACs generated by FBP (blue squares), TRT (black circles), eSMART (gray plus), and TRAINER (red stars) are compared with the ground truth (green circles).

### v2

@@ II-E The Joint Optimization Problem in TRAINER
Where $\mathbf{V}$ denotes the dual variable of $\mathbf{U}$, and $\lambda$ controls the trade-off between data fitting term and the regularization term.
The unconstrained problem can be efficiently solved by the alternating direction method of multipliers (ADMM) algorithm iteratively through the following three sub-problems:
@@ II-F Numerical Optimization Algorithm
The optimization of sub-problem Eq. (17) can be technically challenging. As we can observe, the cost function in Eq. (17) is highly nonlinear and non-convex with respect to the unknown perfusion parametric images. Here, we choose to use the Adam optimizer [30] to minimize the cost function defined in Eq. (17). The Adam optimizer is particularly suitable for this task due to its ability to handle non-smooth objective functions and dynamically adapt learning rates to facilitate faster convergence.
@@
Sub-problem Eq. (19) is a nonlinear least square problem and it has the same form as the conventional supervised training problem, using $\mathbf{U}^{(k+1)}+\mathbf{V}^{(k)}$ as the training labels. Gradient descent has been employed to solve this supervised training problem.
@@ II-F3 Pseudo-code to Implement TRAINER
To implement TRAINER, we first need to initialize $\mathbf{U}$, $\mathbf{V}$, and $\mathbf{\Theta}$, and then optimize these arguments iteratively.
@@
1:  Input: Line integral data $\vec{\mathbf{Y}}$, Initial $\mathbf{CBF}$ and $\mathbf{\mathcal{T}}_{0}$
2:  Output: $\mathbf{U}=\{\mathbf{CBF},\mathbf{\mathcal{T}}_{0}\}$
3:  Initialize: $\mathbf{V}^{(0)},\mathbf{\Theta}^{(0)},\mathbf{Z},\lambda,\eta,\alpha,\varepsilon$
4:  while $k\leq\text{Num Of ADMM}$ and $\mathcal{L}_{\lambda}(\mathbf{U}^{(k)},\mathbf{\Theta}^{(k)},\mathbf{V}^{(k)})\geq\varepsilon$ do
5:     $\mathbf{U}^{(0)}\leftarrow\mathbf{U}^{(k)}$
6:     Update $\mathbf{U}^{(k+1)}$ using Adam optimizer:
7:     for $q=0$ to $Q$ do
8:        Compute $\mathbf{g}_{\mathbf{U}}^{(q)}=\nabla_{\mathbf{U}}\bigg{[}\frac{1}{2}\|\vec{\mathbf{Y}}-\overrightarrow{\bar{\mathbf{Y}}(\mathbf{U})}\|_{\mathbf{W}}^{2}+\frac{\lambda}{2}\|\mathbf{U}-\mathcal{G}_{\mathbf{\Theta}^{(k)}}(\mathbf{Z})+\mathbf{V}^{(k)}\|^{2}\bigg{]}$
9:        $\mathbf{U}^{(q+1)}=\text{AdamUpdate}(\mathbf{U}^{(q)},\mathbf{g}_{\mathbf{U}}^{(q)},\eta)$
10:     end for
11:     $\mathbf{U}^{(k+1)}\leftarrow\mathbf{U}^{(Q)}$
12:     Update $\mathbf{\Theta}^{(k+1)}$ via gradient descent:
13:     for $p=0$ to $P$ do
14:        Compute $\mathbf{g}_{\mathbf{\Theta}}^{(p)}=\nabla_{\mathbf{\Theta}}\|\mathbf{U}^{(k+1)}+\mathbf{V}^{(k)}-\mathcal{G}_{\mathbf{\Theta}}(\mathbf{Z})\|^{2}$
15:        $\mathbf{\Theta}^{(p+1)}=\mathbf{\Theta}^{(p)}-\alpha\mathbf{g}_{\mathbf{\Theta}}^{(p)}$
16:     end for
17:     $\mathbf{\Theta}^{(k+1)}\leftarrow\mathbf{\Theta}^{(P)}$
18:     Update $\mathbf{V}^{(k+1)}$:
19:     $\mathbf{V}^{(k+1)}=\mathbf{V}^{(k)}+\mathbf{U}^{(k+1)}-\mathcal{G}_{\mathbf{\Theta}^{(k+1)}}(\mathbf{Z})$
20:     $k\leftarrow k+1$
@@ II-F4 Network Architecture
The architecture of $\mathcal{G}_{\mathbf{\Theta}}$ is illustrated in Fig. 1 (a). The architecture is designed based on an encoder-decoder structure with skip connections. Parameters $n_{u}[i]$, $n_{d}[i]$, and $n_{s}[i]$ denote the number of filters for upsampling, downsampling, and skip connections at the $i$-th depth, respectively. Parameters $k_{u}[i]$, $k_{d}[i]$, and $k_{s}[i]$ denote the kernel sizes at the $i$-th depth, respectively [31]. Bilinear interpolation is employed for the upsampling process. We use LeakyReLU [32] to introduce the non-linearity. During the iterative training process of network parameters $\mathbf{\Theta}$, we add additive normal noise with a mean of zero and a standard deviation of $\sigma_{p}=\frac{1}{30}$ to input $\mathbf{Z}$ for perturbation. Experimental results show that this method yields superior performance. Additionally, previous research indicates that the optimization process tends to become unstable as the loss decreases and approaches a specific threshold. This instability is characterized by a notable increase in the loss value and the appearance of blurriness in the generated image $\mathcal{G}_{\mathbf{\Theta}}$. To make the training process more stable, we revert the parameters to those from the previous iteration if the loss difference between two consecutive iterations exceeds a specified threshold [31].
Figure 1: (a) The architecture of $\mathcal{G}$. $n_{u}[i]$, $n_{d}[i]$, and $n_{s}[i]$ denote the number of filters for upsampling, downsampling, and skip connections at the $i$-th depth respectively. Parameters $k_{u}[i]$, $k_{d}[i]$, and $k_{s}[i]$ denote the kernel sizes at the $i$-th depth respectively. (b) Conditional image generation. During each iteration, the network parameters $\mathbf{\Theta}$ determine the image $\mathcal{G}_{\mathbf{\Theta}}(\mathbf{Z})$, and the mapping $\mathcal{G}$ is represented as a deep convolutional neural network parameterized by $\mathbf{\Theta}$.
@@ II-G Implementation Details
II-G1 Initialization and Hyperparameter Setting
We first initialized the iterative sequence via the following empirical schemes to obtain the warm initialization of $\mathbf{U}$. The value of CBF for each pixel is initialized as 40 mL/100g/min and $\mathbf{\mathcal{T}}_{0}$ for each pixel is initialized as 5 seconds because such initialization helps the iterative process converge more quickly to the expected solution. $\mathbf{\Theta}$ is initialized as random values uniformly distributed between 0 and 1, while $\mathbf{V}$ is initialized as 0. The variable $\mathbf{Z}$ is initialized within the interval $[0,1/10)$. To optimize the performance of TRAINER, the hyperparameters $\lambda$, $\alpha$, $\eta$ for numerical simulation studies are empirically set to 0.01, 0.01, and 0.001, respectively.
II-G2 Other Remarks
@@ III-A Numerical Simulation Studies with Known Ground Truth
The design of the numerical phantom in this study was based on a digital anthropomorphic perfusion phantom [33]. Regions corresponding to artery, healthy tissue, cerebrospinal fluid, penumbra and ischemic core are defined in the phantom design software toolkit. The residual functions are defined according to programmed cerebral blood flow and mean transit time values. The ground truth TAC of each pixel is then calculated according to the indicator dilution theory. The data acquisition geometry was designed to simulate the scans performed in the animal subject study.
@@
Ideal projection data were generated using a ray-driven implementation of fan-beam forward projection. Projections with quantum noise were then produced by sampling from a Poisson distribution for each projection ray of the ideal data. For CBCT perfusion data acquisitions, an entrance photon number of $I_{0}=1\times 10^{6}$ photons per ray was used to simulate the full exposure level. Entrance photon number per ray $I_{0}=1\times 10^{5},5\times 10^{4},1\times 10^{4}$ were used to simulate lower exposure levels respectively.
@@ III-B Baseline Methods for Performance Comparison
Four baseline methods are compared: the Singular Value Decomposition (SVD)-based with Tikhonov regularization algorithms [34], and Tensor Total-Variation Regularization (TTV) [35], temporal recovery technique (TRT) [10], and eSMART-RECON [6]. For the (SVD)-based with Tikhonov regularization algorithms, threshold value of 0.15 (15% of the maximum singular value) is empirically chosen to achieve optimal performance. These methods are widely recognized as regularized deconvolution techniques for CT perfusion imaging and are extensively incorporated into commercial medical software [36]. TTV [35] applies total variation regularization to the impulse response function, integrating spatial vascular structure correlation and temporal blood signal continuity. This algorithm converges rapidly and exhibits low computational complexity. In the implementation of the TTV algorithm in this study, both the regularization parameters for the temporal and spatial dimensions, denoted as $\gamma$, were set to 0.001, with a total of 60 iterations performed. TRT algorithm introduces a novel temporal recovery technique designed to restore the time attenuation curves in C-arm CBCT perfusion imaging, addressing the limitations of existing C-arm CBCT systems in terms of data acquisition speed and temporal sampling density. In the implementation of the TRT algorithm in this study, the correction factor was set to 0.5. eSMART improves the temporal resolution of C-arm CBCT via incorporating the prior knowledge of periodicity of the limited-view artifacts in the C-arm CBCT perfusion data acquisition scheme. For each time point, the eSMART algorithm selects the 10 adjacent frames of projection data as input, and performs 15 iterations. A relaxation factor of 0.06 is introduced during the iteration process to optimize both iteration efficiency and the quality of the reconstructed images. To improve image quality and quantitative accuracy at lower exposure levels, noise reduction in the time-resolved images reconstructed by each baseline method was achieved through spatial and temporal domain filtering. Specifically, spatial filtering was performed using a Gaussian filter with a standard deviation of 1 and a kernel size of $5\times 5$, while temporal filtering was conducted with a Gaussian filter having a standard deviation of 1 and a kernel size of $3\times 1$. Perfusion parametric images of each baseline method were then generated using the corresponding denoised time-resolved images.
@@ III-C Metrics for Quantitative Performance Assessment
The study employs relative Root Mean Square Error (rRMSE) and Peak Signal-to-Noise Ratio (PSNR) as evaluation metrics.
$\mathrm{rRMSE}=\frac{\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-y_{i}^{\text{truth}}\right)^{2}}}{\max y^{\text{truth}}}\times 100\%,$ (22)
where $y$ denotes the image being assessed, $y^{\text{truth}}$ denotes the ground truth image, $\max y^{\text{truth}}$ denotes the maximum pixel value of the ground truth image, and the subscript $i$ denotes the $i$-th voxel index of the image.
@@ IV-B Qualitative Assessment of Perfusion Parametric Images
Quantitative perfusion images reconstructed at the gantry rotation speed of $dt=8$ sec and four exposure levels are shown in Fig. 3 and Fig. 4. For CBF images, at the full exposure level, compared with the ground truth, FBP+Tikh and FBP+TTV underestimate healthy tissue while overestimate diseased tissues (penumbra and ischemic core). TRT+TTV underestimates but eSMART+TTV overestimates healthy tissue. In addition, both TRT+TTV and eSMART+TTV overestimate diseased tissues. In contrast, TRAINER can accurately reconstruct CBF values for all types of tissues. When the exposure level is getting lower, penumbra and ischemic core are not distinguished in baseline methods. Error and noise level of CBF images reconstructed by baseline methods are getting larger. For MTT images, at the full exposure level, compared with the ground truth, baseline methods underestimate all tissue types. Penumbra and ischemic core are not distinguished in baseline methods even at the full exposure level. TRAINER accurately estimate the value of all types of tissues. When the exposure level is getting lower, error and noise level of MTT images reconstructed by baseline methods are getting larger. Results shown in Fig. 3 and Fig. 4 demonstrated that TRAINER can accurately reconstruct quantitative perfusion images at gantry rotation speed of $dt=8$ sec even at the lowest exposure level for all tissue types while other methods cannot.
@@
To validate the effectiveness of the proposed TRAINER and also explain why other baseline methods cannot obtain accurate quantitative perfusion images, we presented the time attenuation curves (TACs) of the reconstructed time-resolved images using TRAINER and compared them to those obtained from the baseline methods.
As shown in Fig. 5, TRAINER consistently generates TACs closer to the ground truth (Truth) across three different tissue types: healthy tissue, penumbra, and ischemic core. Specifically, in the healthy tissue region, TACs obtained using TRAINER are nearly consistent with Truth, demonstrating a significant improvement on reconstruction accuracy over the baseline methods. In the penumbra region, although all methods show some fluctuations, TACs generated by TRAINER remains closer to Truth. In the ischemic core region, TRAINER yields smoother and more accurate TACs compared to the baseline methods, further highlighting the superior quantitative accuracy of TRAINER. These results further confirm the efficacy of TRAINER, demonstrating its superior performance in reconstructing TACs using data acquired by slow rotation speed systems.
Figure 5: Comparison of time attenuation curves for different tissue types. The plots illustrate the time attenuation curves for healthy tissue (top), penumbra (middle), and ischemic core (bottom). The TACs generated by FBP (blue squares), TRT (black circles), eSMART (gray plus), and TRAINER (red stars) are compared with the ground truth (green circles).